OpenStudy (anonymous):

solve using elimintion : 3x + 2y = 13 x + 2y = 7

OpenStudy (anonymous):

So first, you can solve the second equation for x rather easily. Simply subtract 2y from both sides and you have x isolated. Then, substitute x = 7-2y into the first equation. Then you will have an equation that is JUST in terms of y and you can solve for y. Once you find y, plug in it's value into either of the original equations and solve for x.

OpenStudy (anonymous):

can you put that in numbers for me ? cos im a slow person .

OpenStudy (anonymous):

Certainly. x = 7-2y This implies that 3(7-2y) + 2y = 13

OpenStudy (anonymous):

So, 21 - 4y + 2y = 13 Now you can solve for y. Once you have y, x is a synch.

OpenStudy (anonymous):

Actually, the method I used was back substitution. If you want to solve using elimination, you have to add multiples of equations to cancel out variables. If you call the first equation A and the second equation B, then to get rid of x you would do the following: A - 3B Which would yield (3x + 2y) - 3(x+2y) = 13 - 3*7

OpenStudy (anonymous):

In that case, you "eliminate" x and then just have an equation in terms of y. Perhaps it's easier to eliminate y first. In that case, you would simply do the following A-B Which would yield (3x+2y) - (x+2y) = 13 - 7 In THIS case, you will end out with an equation that ONLY has x in it. Therefore, you will be able to solve for x.

OpenStudy (anonymous):

So, the idea with elimination is that you can add and subtract full equations like I showed previously. You always want to add or subtract the equations so that you "eliminate" one variable to make the equation easy to solve.